OpenStudy (anonymous):

PLEASE HELP ASAP When he was 35, Keene began investing $250 per month in various securities for his retirement savings. His investments averaged a 7.5% annual rate of return until he retired at age 60. What was the value of Keene’s retirement savings when he retired? Assume monthly compounding of interest. $259,315.22 $336,861.36 $877.26 $219,315.22

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathmate):

Are you familiar with the amortization formula?

OpenStudy (anonymous):

no

OpenStudy (mathmate):

If someone puts in each month $A, and for n months at annual interest of i%, then the final amount is P. where \( P = \frac{A(R^n-1)}{(R-1)} \) where P=final amount R=1+interest rate per month (1+7.5%/12 = 1.00625) n=number of months before he retires = 25*12

OpenStudy (mathmate):

Did you get the answer?